# Algorithmic Decision-Making under Agents with Persistent Improvement

### v1

@@ Appendix D Illustration of Table 1
Appendix E Illustration of Thm. 5.1
Refer to captionRefer to caption
Figure 8: Illustration of Thm. 5.1: the left figure shows $\widetilde{U}-U$ as functions of $x_{0}$ under different $P$ when $\theta=0.995,r=0.05$; the right plot shows threshold $\widehat{P}$ under different pairs of $(\theta,r)$.
@@
Thm. 5.1 identifies conditions under which manipulation (or improvement) is preferred by individuals over the other. As mentioned in Section 5, the specific values of $\widehat{P}$, $\widehat{x}$, $\widehat{x_{1}}$, $\widehat{x_{2}}$ in Thm. 5.1 depend on $\theta$, $r$, and we can empirically find $\widehat{P}$, $\widehat{x}$, $\widehat{x_{1}}$, $\widehat{x_{2}}$ and verify the theorem, as illustrated in Figure 8 and Table 2. Specifically, the left plot in Figure 8 shows $\widetilde{U}-U$ as functions of initial similarity $x_{0}$ under different detection probability $P$. Because individuals only prefer to manipulate if $\widetilde{U}-U>0$, the plot shows the values of $\widehat{P}$, $\widehat{x}$, $\widehat{x_{1}}$, $\widehat{x_{2}}$ in Thm. 5.1. The right plot shows threshold $\widehat{P}$ under different pairs of $(\theta,r)$, and it shows that $\widehat{P}$ increases as $r$ increases. Table 2 shows ranges $(\widehat{x_{1}},\widehat{x_{2}})$ of initial similarity $x_{0}$ under different detection probability $P$, acceptance threshold $\theta$, and discounting factor $r$.
Appendix F Additional Experiments
@@
Just as Sec. 7 mentions, we acquire the exam score data Kimmons (2012), preprocess the data and fit beta distributions for both males and females. The fitted distribution and real distribution are shown in Fig. 9.
Refer to captionRefer to caption
@@
Just as Sec. 7 mentions, we fit beta distributions for FICO Score Hardt et al. (2016b), and obtain four distributions for different racial groups as shown in Fig. 10.
Additional Results for FICO Data
Besides Caucasian and African American mentioned in Sec. 7, for Asians and Hispanic, we also compute the optimal decision threshold and corresponding total improvement under different $r$. As shown in Fig. 11, $\theta^{*}$ always decreases with $r$ and the total amount of improvement decreases. If comparing Asians and Hispanics, we observe that Hispanics have lower thresholds but larger improvements. For settings with both manipulation and improvement (Fig. 11), it seems that a larger (resp. smaller) proportion of Asians tend to manipulate than African Americans under $\theta^{*}$. More importantly, the optimal thresholds reveal larger amounts of improvement for Hispanics, suggesting that the decision-maker’s policy in Sec. 4 is beneficial for the disadvantaged group.
@@
Appendix G Estimating the discounting factor $r$ in Sec.4
@@
Appendix H Proofs
H.1 Proof Details of Thm. 3.1
@@ Appendix H Proofs
H.2 Proof Details of Thm. 4.1 and Corollary 4.2
H.2.1 Proof of Thm. 4.1
@@
H.2.2 Proofs of why Uniform distribution has a unique maximized $\theta^{*}$
@@ 1. Situation when P = 0
Firstly, we prove when $x_{0}\rightarrow\theta$, $U^{*}<\widetilde{U}$: when $x_{0}\rightarrow\theta$, we know $k^{*}\rightarrow 0$ since individuals invest an arbitrarily small effort to immediately qualified. However, according to Sec. H.1, $k^{*}$ should let $\frac{\partial^{2}U}{\partial k^{2}}<0$. This inequality will give us the bound of $k^{*}$: $\ln(k^{*}+1)>\frac{-G(x_{0})\cdot ln(1+r)}{3}$. With this bound, we can plug $k^{*}$ into (19), and know $\frac{ln(1+r)}{ln(k^{*}+1)}\rightarrow+\infty$, and $(1+r)^{\frac{G(x_{0})}{ln(k^{*}+1)}}$ is larger than a constant because of the bound. Therefore, $\frac{\partial U^{*}}{\partial x_{0}}\geq\frac{\partial G}{\partial x_{0}}\cdot+\infty$. Then according to (20), when $x_{0}\rightarrow\theta$, $\frac{\partial\widetilde{U}}{\partial x_{0}}<\frac{\partial G}{\partial x_{0}}\cdot\frac{ln(1+r)}{ln2}+1$. Since $\frac{\partial G}{\partial x_{0}}$ is always positive, when $x_{0}\rightarrow\theta$, we prove that $\frac{\partial U^{*}}{\partial x_{0}}>\frac{\partial\widetilde{U}}{\partial x_{0}}$. Meanwhile, when $x_{0}=\theta$, $U^{*}=\widetilde{U}=1$. This means when $x_{0}\rightarrow\theta$, $U^{*}<\widetilde{U}$.
@@
Thus. there must be an intersection between $\widetilde{U}$ and $U^{*}$. Then noticing that if we increase $\theta$, $\widetilde{U}$ is always decreasing to converge to function $y=x-1$, while $U^{*}\geq 0$ always holds. This suggests when $\theta$ is sufficiently close to 1, we can guarantee the first intersection of $U^{*}$ and $\widetilde{U}$ occurs arbitrarily close to 1, meaning this first intersection is the only intersection.
@@ 2. Situation when P $>$ 0
H.4 Proof Details of Thm. 6.2
@@
Proof
According to Thm. 6.1, $k_{u}^{2}=\|\widetilde{d}\|^{2}\cdot x_{0}^{2}$ and $\|\widetilde{d}\|^{2}=k^{2}+(1-k)^{2}-2k(1-k)x_{0}$. We can get following expression:
@@
Proof. If we let $z=\ln(k+1)$ be $z$ and recall that $C(\theta,x_{0},r)=-\ln\left(\sqrt{\frac{(\theta)^{-2}-1}{(x_{0})^{-2}-1}}\right)\cdot\ln(1+r)$, we would have $U=e^{\frac{-C}{z}}-e^{z}+1$.
@@
Proof of Thm. 6.2
According to Lemma H.1, when $\bar{k}\in[0,\widehat{k})$, $k_{u}>k$, so the convergence speed of the individual to $d^{*}$ under forgetting mechanism will be faster than the convergence speed of the individual to $d$ without forgetting mechanism, so that the reward under forgetting mechanism is discounting less. Meanwhile, according to Lemma H.2, there exists a $r$ where $U(\bar{k},\theta,r,x_{0})>0$. Combine them together, $\widehat{U}(\bar{k},\theta,r,x_{0})>U(\bar{k},\theta,r,x_{0})>0$ and Thm. 6.2 is proved.
H.5 Proof of Thm. B.1
Assume $k_{t}=\frac{k}{t+1}$ when $t\geq 0$. From (1) and similar to Dean and Morgenstern (2022), we know $(q_{t+1}^{T}\cdot d)^{-2}-1=\frac{(q_{t}^{T}\cdot d)^{-2}-1}{k_{t}+1}^{2}$. This will lead to $(q_{t}^{T}\cdot d)^{-2}-1=\prod_{i=0}^{t-1}(\frac{k}{i+1}+1)^{-2}((q_{0}^{T}\cdot d)^{-2}-1)$.

### v2

@@ Appendix B Illustration of Table 1
Appendix C Illustration of Thm. 6.1
Figure 9: Illustration of Thm. 6.1: the left figure shows $\widetilde{U}-U$ as functions of $x_{0}$ under different $P$ when $\theta=0.995,r=0.05$; the right plot shows threshold $\widehat{P}$ under different pairs of $(\theta,r)$.
@@
Thm. 6.1 identifies conditions under which manipulation (or improvement) is preferred by individuals over the other. As mentioned in Section 6, the specific values of $\widehat{P}$, $\widehat{x}$, $\widehat{x_{1}}$, $\widehat{x_{2}}$ in Thm. 6.1 depend on $\theta$, $r$, and we can empirically find $\widehat{P}$, $\widehat{x}$, $\widehat{x_{1}}$, $\widehat{x_{2}}$ and verify the theorem, as illustrated in Figure 9 and Table 2. Specifically, the left plot in Figure 9 shows $\widetilde{U}-U$ as functions of initial similarity $x_{0}$ under different detection probability $P$. Because individuals only prefer to manipulate if $\widetilde{U}-U>0$, the plot shows the values of $\widehat{P}$, $\widehat{x}$, $\widehat{x_{1}}$, $\widehat{x_{2}}$ in Thm. 6.1. The right plot shows threshold $\widehat{P}$ under different pairs of $(\theta,r)$, and it shows that $\widehat{P}$ increases as $r$ increases. Table 2 shows ranges $(\widehat{x_{1}},\widehat{x_{2}})$ of initial similarity $x_{0}$ under different detection probability $P$, acceptance threshold $\theta$, and discounting factor $r$.
Appendix D Additional Experiments
@@
Just as Sec. 8 mentions, we acquire the exam score data (Kimmons 2012), preprocess the data and fit beta distributions for both males and females. The fitted distribution and real distribution are shown in Fig. 10.
@@
Just as Sec. 8 mentions, we fit beta distributions for FICO Score (Hardt et al. 2016b), and obtain four distributions for different racial groups as shown in Fig. 11.
@@
Appendix E Estimating the discounting factor $r$ in Sec.5
@@
Appendix F Proofs
F.1 Proof Details of Thm. 4.1
@@ Appendix F Proofs
F.2 Proof Details of Thm. 5.1 and Corollary 5.2
Proof of Thm. 5.1
@@
Proofs of why Uniform distribution has a unique maximized $\theta^{*}$
@@ 1. Situation when P = 0:
Firstly, we prove when $x_{0}\rightarrow\theta$, $U^{*}<\widetilde{U}$: when $x_{0}\rightarrow\theta$, we know $k^{*}\rightarrow 0$ since individuals invest an arbitrarily small effort to immediately qualified. However, according to Sec. F.1, $k^{*}$ should let $\frac{\partial^{2}U}{\partial k^{2}}<0$. This inequality will give us the bound of $k^{*}$: $\ln(k^{*}+1)>\frac{-G(x_{0})\cdot ln(1+r)}{3}$. With this bound, we can plug $k^{*}$ into (20), and know $\frac{ln(1+r)}{ln(k^{*}+1)}\rightarrow+\infty$, and $(1+r)^{\frac{G(x_{0})}{ln(k^{*}+1)}}$ is larger than a constant because of the bound. Therefore, $\frac{\partial U^{*}}{\partial x_{0}}\geq\frac{\partial G}{\partial x_{0}}\cdot+\infty$. Then according to (21), when $x_{0}\rightarrow\theta$, $\frac{\partial\widetilde{U}}{\partial x_{0}}<\frac{\partial G}{\partial x_{0}}\cdot\frac{ln(1+r)}{ln2}+1$. Since $\frac{\partial G}{\partial x_{0}}$ is always positive, when $x_{0}\rightarrow\theta$, we prove that $\frac{\partial U^{*}}{\partial x_{0}}>\frac{\partial\widetilde{U}}{\partial x_{0}}$. Meanwhile, when $x_{0}=\theta$, $U^{*}=\widetilde{U}=1$. This means when $x_{0}\rightarrow\theta$, $U^{*}<\widetilde{U}$.
@@
Thus. there must be an intersection between $\widetilde{U}$ and $U^{*}$. Then noticing that if we increase $\theta$, $\widetilde{U}$ is always decreasing to converge to function $y=x-1$, while $U^{*}\geq 0$ always holds. This suggests when $\theta$ is sufficiently close to 1, i.e., there exists a $\bar{\theta}$ and when $\theta>\bar{\theta}$, we can guarantee the first intersection of $U^{*}$ and $\widetilde{U}$ occurs arbitrarily close to 1, meaning this first intersection is the only intersection.
@@ 2. Situation when P $>$ 0:
F.4 Proof Details of Thm. 7.1
Thm. 7.1 can be proved by the inequality $\|\widetilde{d}\|q_{t}^{T}d>\|\widetilde{d}\|q_{0}^{T}d>\|\widetilde{d}\|q_{0}^{T}dq_{t}^{T}d^{*}$. This means we can just let $k_{u}=\|\widetilde{d}\|q_{0}^{T}d=\|\widetilde{d}\|x_{0}$ and directly apply Lemma 3.2.
F.5 Proof Details of Thm. 7.2
@@
Proof.
According to Thm. 7.1, $k_{u}^{2}=\|\widetilde{d}\|^{2}\cdot x_{0}^{2}$ and $\|\widetilde{d}\|^{2}=k^{2}+(1-k)^{2}-2k(1-k)x_{0}$. We can get following expression:
@@
Proof.
If we let $z=\ln(k+1)$ be $z$ and recall that $C(\theta,x_{0},r)=-\ln\left(\sqrt{\frac{(\theta)^{-2}-1}{(x_{0})^{-2}-1}}\right)\cdot\ln(1+r)$, we would have $U=e^{\frac{-C}{z}}-e^{z}+1$.
@@
Proof of Thm. 7.2.
According to Lemma F.1, when $\bar{k}\in[0,\widehat{k})$, $k_{u}>k$, so the convergence speed of the individual to $d^{*}$ under forgetting mechanism will be faster than the convergence speed of the individual to $d$ without forgetting mechanism, so that the reward under forgetting mechanism is discounting less. Meanwhile, according to Lemma F.2, there exists a $r$ where $U(\bar{k},\theta,r,x_{0})>0$. Combine them together, $\widehat{U}(\bar{k},\theta,r,x_{0})>U(\bar{k},\theta,r,x_{0})>0$ and Thm. 7.2 is proved.
F.6 Proof of Thm. A.1
Assume $k_{t}=\frac{k}{t+1}$ when $t\geq 0$. From (1) and similar to (Dean and Morgenstern 2022), we know $(q_{t+1}^{T}\cdot d)^{-2}-1=\frac{(q_{t}^{T}\cdot d)^{-2}-1}{k_{t}+1}^{2}$. This will lead to $(q_{t}^{T}\cdot d)^{-2}-1=\prod_{i=0}^{t-1}(\frac{k}{i+1}+1)^{-2}((q_{0}^{T}\cdot d)^{-2}-1)$.